how do you find the slope for y - 3= 2/7(x+4)??
Found 2 solutions by persian52, Alan3354:
Answer by persian52(161)

y-3=(2)/(7)*(x+4)
Multiply 2 by each term inside the parentheses.
y-3=((2x+8))/(7)
Remove the parentheses around the expression 2x+8.
y-3=(2x+8)/(7)
Divide each term in the numerator by the denominator.
y-3=(2x)/(7)+(8)/(7)
Since -3 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3 to both sides.
y=3+(2x)/(7)+(8)/(7)
Simplify the right-hand side of the equation.
y=(2x+29)/(7)
Divide each term in the numerator by the denominator.
y=(2x)/(7)+(29)/(7)
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=(2)/(7) and b=(29)/(7).
Slope: m=(2)/(7)
Y-intercept: b=(29)/(7)

Answer by Alan3354(69443)

It's the coefficient of x when the coefficient of y is 1.
m = 2/7
The other terms have no effect on the slope.